OpenStudy (anonymous):

A population has a mean of 50 and a standard deviation of 16. If a random sample of 49 is taken, what is the probability that the sample mean is each of the following? a. Greater than 57

OpenStudy (anonymous):

Impossible to answer without information about the distribution.

OpenStudy (zarkon):

I'm sure they are supposed to use the central limit theorem here.

OpenStudy (anonymous):

^what he said

OpenStudy (zarkon):

\[P(\overline{x}>57)=P\left(\frac{\overline{x}-50}{16/\sqrt{49}}>\frac{57-50}{16/\sqrt{49}}\right)\] then by clt \[=P\left(Z>3.0625\right)=1-P(Z\le3.0625)=\cdots\]